### v1

@@ Abstract
As a promising avenue for future research, a satellite platform with maneuvering capabilities can potentially explore the gravitational redshift of single photons using single-photon interference rideout2012fundamental ; lu2022micius . For example, Fig. 1a illustrates the measurement of single-photon gravitational redshift using a Mach-Zehnder interferometer (MZI). Clearly, on the two paths of the MZI, the blue dashed parts pass through the same gravitational field zych2011quantum ; zych2012general . Meanwhile, the yellow dotted and green dash-dotted portions traverse different gravitational fields, introducing gravitational-induced phase shifts. However, the phase noise introduced by the ground-to-satellite channel in the two arms of the MZI swamps the gravitational redshift phase, rendering the experiment impossible. To overcome this challenge, a Franson interferometer can be utilized in Fig.1b, consisting of two independent unbalanced Michelson interferometers (UMIs) and a shared free-space channel, equivalent to the MZI rideout2012fundamental . The shared free-space channel enables the cancellation of atmospheric noise as common-mode noise. Assuming a satellite orbiting at an altitude of $h$ and an unbalanced arm of the interferometer of $\delta l$, the redshift phase can be expressed as
@@
Figure 1: a, In the MZI, single photons coherently travel along the left blue dashed and green dash-dotted paths or the right blue dashed and yellow dotted paths. Since the two blue dashed paths traverse the same gravitational field and cancel each other out, the gravitational redshift effect will introduce the wavelength difference of single photons between the green dash-dotted and yellow dotted paths, resulting in a phase shift in single photon interference zych2011quantum ; zych2012general . b, To simplify the setup, two blue dashed paths are combined into one, while the yellow dotted and green dash-dotted paths become the unbalanced arm lengths of two UMI in a Franson interferometer rideout2012fundamental .
@@
In our scheme, single photons are generated from a QDSPS and adjusted to state $\frac{1}{\sqrt{2}}(|H\rangle+|V\rangle)$ using a polarization beam splitter (PBS) and a half waveplate (HWP), as shown in Fig. 2a. The single photons are then fed into the UMI, which transforms the state to $|\Psi_{t}\rangle=\frac{1}{\sqrt{2}}(|H\rangle|S\rangle+e^{i\varphi_{t}}|V\rangle|L\rangle)$. Here, $|L\rangle$ and $|S\rangle$ denote the states passing through the long and short arms of the UMI, respectively, and $\varphi_{t}$ is a constant internal phase of the UMI at the transmitter. We placed the UMI in a vacuum chamber to achieve high-precision phase measurement and recorded the air pressure values using a computer, which allowed us to isolate the effects of air pressure fluctuations on our measurements. The delay time between $|L\rangle$ and $|S\rangle$ is $\delta t\approx 4$ ns, corresponding to the UMI’s arm difference of $\delta l=c\delta t\approx 1.2$ m ($c$ is the speed of light in vacuum). The single photons were then coupled into single-mode fiber and guided to a launch telescope.
@@
After traveling through the free-space channel in the Shanghai metropolitan area, the single photons were captured by the receiving system, as shown in Fig. 2b. We used adjustable waveplates to introduce a variable phase $\varphi$ between $|H\rangle$ and $|V\rangle$ state for interference visibility measuring. After passing through the second UMI, the state changes to $|\Psi_{r}\rangle=\frac{1}{\sqrt{2}}|L\rangle|S\rangle(|H\rangle+e^{i(\varphi_{t}+\varphi+\varphi_{r})}|V\rangle)$, where $\varphi_{r}$ is the phase induced by the UMI at the receiver. Finally, we measured the photon under the Pauli operator $\sigma_{x}$ using an HWP and a PBS, collected photons with multi-mode fibers, with a mode field diameter of 105 µm, supporting up to approximately 40000 different guided modes, and detected using single-photon avalanche detectors (SPAD).
@@
Figure 2: Illustration of experimental setup. a, The QDSPS and UMI generate single photons in a time-bin qubit, which is crucial for achieving high-precision measurements. b, The optical receiving system is mounted on a high-load pitch and yaw platform for coarse pointing. The telescope has an aperture of 127 mm and a beam reduction ratio of 37.5. The received 671-nm-wavelength laser is separated by a dichroic mirror (DM) and imaged by a camera for fine tracking of the fast steering mirror. A motorized rotation stage controls the three waveplates to achieve dynamic modulation of the interference phase. We also use a 10-nm-wide interference filter (IF) to reduce the dark count. c, The scheme and photo of the UMI. All optical elements adhere to an optical substrate that measures 160 mm in length, 80 mm in width, and 30 mm in height. To keep the UMI compact, we adopt a round-trip optical design and append imaging systems in the long arm of the UMI. The red arrow represents the direction of the beam propagation in a photo of the UMI, and the interferometer is put in a vacuum chamber with a temperature controlling system. QWP, quarter waveplate; mFPC, manual fiber polarization controller; FSM, fast steering mirror.
@@
Due to the unbalanced arm length in UMI, its interference performance is significantly affected by atmospheric turbulence, including incident angle jitter and wavefront distortion. This effect can cause reduced interference visibility and degraded phase stability. The spatial filter, such as single-mode fiber coupling, is generally effective, while it entails much more significant attenuation. An alternative solution is to incorporate an imaging system into the UMI. This approach can address the above issues while maintaining good multi-mode interference performance jin2018demonstration ; jin2019genuine . The imaging system comprises two lenses and reduces mode-dependent path length, as illustrated in FIG. 2c. Thus, it ensures high interference visibility and phase stability despite temporal and spatial photon distortion between different arms. Since photons in the long arm pass through the imaging system twice, any spatial mode transformation becomes a nearly identical matrix that similarly affects photons in both arms. To evaluate the efficacy of this imaging system, we introduced an angular change of 1.6 mrad to the incident light of the UMI. Consequently, there was a variation of 0.04 mrad and 0.06 mm in the overlap of the beam’s direction and position, respectively, for the beam passing through the long and short arms, which are much smaller than the spot diameter of 2 mm and the divergence angle of 0.55 mrad.
@@
We measured the multi-mode interference visibility of single-photon interference in an 8.4-km horizontal atmosphere channel located in Shanghai. Here, we use $\frac{C_{1}-C_{2}}{C_{1}+C_{2}}$ to measure interference visibility, where $C_{1}$ and $C_{2}$ represent the single-photon counts of two SPADs. By adjusting $\varphi$, $C_{1}$ can be optimized to reach its maximum value, at which point $\frac{C_{1}-C_{2}}{C_{1}+C_{2}}$ becomes the interference visibility $\mathcal{V}$. In FIG. 3a, three wave plates in the receiving system introduce dynamic phase $\varphi$ for multi-mode interference. The measured multi-mode interference visibility $\mathcal{V}=0.863\pm 0.004$, indicates the effectiveness of the imaging system against atmospheric turbulence. By directly measuring the phase at a frequency of 0.1 Hz under fixed phase $\varphi$, we achieved a phase measurement precision of 35.8 mrad in FIG. 3b. However, a long-term linear phase drift was observed and found to be $0.117\pm 0.006$ mrad/s, which is shown as a blue dashed line. Our previous noise analysis revealed that the temperature drift of the UMI causes long-term phase drift, calculated as 0.137 mrad/s/K. To maintain temperature stability, the air conditioner was set to a constant nighttime temperature during the experiments. Without the long-term phase drift, the short-term stability of phase measurement precision 16.2 mrad is shown as a gray area. Through noise analysis, we identified the inconsistency of SPADs and shot noise as the primary sources of noise. The shot-noise-induced phase noise was calculated to be 4.3 mrad. The inconsistency of SPADs would inevitably introduce phase noise under various free-space channel attenuation, estimated as 15.6 mrad. More details are provided in the Methods.

### v2

@@ Abstract
As a promising avenue for future research, a satellite platform with maneuvering capabilities can potentially explore the gravitational redshift of single photons using single-photon interference rideout2012fundamental ; lu2022micius . For example, Fig. 1(a) illustrates the measurement of single-photon gravitational redshift using a Mach-Zehnder interferometer (MZI). Clearly, on the two paths of the MZI, the blue dashed parts pass through the same gravitational field zych2011quantum ; zych2012general . Meanwhile, the yellow dotted and green dash-dotted portions traverse different gravitational fields, introducing gravitational-induced phase shifts. However, the phase noise introduced by the ground-to-satellite channel in the two arms of the MZI swamps the gravitational redshift phase, rendering the experiment impossible. To overcome this challenge, a Franson interferometer can be utilized in Fig.1(b), consisting of two independent unbalanced Michelson interferometers (UMIs) and a shared free-space channel, equivalent to the MZI rideout2012fundamental . The shared free-space channel enables the cancellation of atmospheric noise as common-mode noise. Assuming a satellite orbiting at an altitude of $h$ and an unbalanced arm of the interferometer of $\delta l$, the redshift phase can be expressed as
@@
Figure 1: (a), In the MZI, single photons coherently travel along the left blue dashed and green dash-dotted paths or the right blue dashed and yellow dotted paths. Since the two blue dashed paths traverse the same gravitational field and cancel each other out, the gravitational redshift effect will introduce the wavelength difference of single photons between the green dash-dotted and yellow dotted paths, resulting in a phase shift in single photon interference zych2011quantum ; zych2012general . (b), To simplify the setup, two blue dashed paths are combined into one, while the yellow dotted and green dash-dotted paths become the unbalanced arm lengths of two UMI in a Franson interferometer rideout2012fundamental .
@@
In our scheme, single photons are generated from a QDSPS and adjusted to state $\frac{1}{\sqrt{2}}(|H\rangle+|V\rangle)$ using a polarization beam splitter (PBS) and a half waveplate (HWP), as shown in Fig. 2(a). The single photons are then fed into the UMI, which transforms the state to $|\Psi_{t}\rangle=\frac{1}{\sqrt{2}}(|H\rangle|S\rangle+e^{i\varphi_{t}}|V\rangle|L\rangle)$. Here, $|L\rangle$ and $|S\rangle$ denote the states passing through the long and short arms of the UMI, respectively, and $\varphi_{t}$ is a constant internal phase of the UMI at the transmitter. We placed the UMI in a vacuum chamber to achieve high-precision phase measurement and recorded the air pressure values using a computer, which allowed us to isolate the effects of air pressure fluctuations on our measurements. The delay time between $|L\rangle$ and $|S\rangle$ is $\delta t\approx 4$ ns, corresponding to the UMI’s arm difference of $\delta l=c\delta t\approx 1.2$ m ($c$ is the speed of light in vacuum). The single photons were then coupled into single-mode fiber and guided to a launch telescope.
@@
After traveling through the free-space channel in the Shanghai metropolitan area, the single photons were captured by the receiving system, as shown in Fig. 2(b). We used adjustable waveplates to introduce a variable phase $\varphi$ between $|H\rangle$ and $|V\rangle$ state for interference visibility measuring. After passing through the second UMI, the state changes to $|\Psi_{r}\rangle=\frac{1}{\sqrt{2}}|L\rangle|S\rangle(|H\rangle+e^{i(\varphi_{t}+\varphi+\varphi_{r})}|V\rangle)$, where $\varphi_{r}$ is the phase induced by the UMI at the receiver. Finally, we measured the photon under the Pauli operator $\sigma_{x}$ using an HWP and a PBS, collected photons with multi-mode fibers, with a mode field diameter of 105 µm, supporting up to approximately 40000 different guided modes, and detected using single-photon avalanche detectors (SPAD).
@@
Figure 2: Illustration of experimental setup. (a), The QDSPS and UMI generate single photons in a time-bin qubit, which is crucial for achieving high-precision measurements. b, The optical receiving system is mounted on a high-load pitch and yaw platform for coarse pointing. The telescope has an aperture of 127 mm and a beam reduction ratio of 37.5. The received 671-nm-wavelength laser is separated by a dichroic mirror (DM) and imaged by a camera for fine tracking of the fast steering mirror. A motorized rotation stage controls the three waveplates to achieve dynamic modulation of the interference phase. We also use a 10-nm-wide interference filter (IF) to reduce the dark count. (c), The scheme and photo of the UMI. All optical elements adhere to an optical substrate that measures 160 mm in length, 80 mm in width, and 30 mm in height. To keep the UMI compact, we adopt a round-trip optical design and append imaging systems in the long arm of the UMI. The red arrow represents the direction of the beam propagation in a photo of the UMI, and the interferometer is put in a vacuum chamber with a temperature controlling system. QWP, quarter waveplate; mFPC, manual fiber polarization controller; FSM, fast steering mirror.
@@
Due to the unbalanced arm length in UMI, its interference performance is significantly affected by atmospheric turbulence, including incident angle jitter and wavefront distortion. This effect can cause reduced interference visibility and degraded phase stability. The spatial filter, such as single-mode fiber coupling, is generally effective, while it entails much more significant attenuation. An alternative solution is to incorporate an imaging system into the UMI. This approach can address the above issues while maintaining good multi-mode interference performance jin2018demonstration ; jin2019genuine . The imaging system comprises two lenses and reduces mode-dependent path length, as illustrated in Fig. 2(c). Thus, it ensures high interference visibility and phase stability despite temporal and spatial photon distortion between different arms. Since photons in the long arm pass through the imaging system twice, any spatial mode transformation becomes a nearly identical matrix that similarly affects photons in both arms. To evaluate the efficacy of this imaging system, we introduced an angular change of 1.6 mrad to the incident light of the UMI. Consequently, there was a variation of 0.04 mrad and 0.06 mm in the overlap of the beam’s direction and position, respectively, for the beam passing through the long and short arms, which are much smaller than the spot diameter of 2 mm and the divergence angle of 0.55 mrad.
@@
We measured the multi-mode interference visibility of single-photon interference in an 8.4-km horizontal atmosphere channel located in Shanghai. Here, we use $\frac{C_{1}-C_{2}}{C_{1}+C_{2}}$ to measure interference visibility, where $C_{1}$ and $C_{2}$ represent the single-photon counts of two SPADs. By adjusting $\varphi$, $C_{1}$ can be optimized to reach its maximum value, at which point $\frac{C_{1}-C_{2}}{C_{1}+C_{2}}$ becomes the interference visibility $\mathcal{V}$. In Fig. 3(a), three wave plates in the receiving system introduce dynamic phase $\varphi$ for multi-mode interference. The measured multi-mode interference visibility $\mathcal{V}=0.863\pm 0.004$, indicates the effectiveness of the imaging system against atmospheric turbulence. By directly measuring the phase at a frequency of 0.1 Hz under fixed phase $\varphi$, we achieved a phase measurement precision of 35.8 mrad in Fig. 3(b). However, a long-term linear phase drift was observed and found to be $0.117\pm 0.006$ mrad/s, which is shown as a blue dashed line. Our previous noise analysis revealed that the temperature drift of the UMI causes long-term phase drift, calculated as 0.137 mrad/s/K. To maintain temperature stability, the air conditioner was set to a constant nighttime temperature during the experiments. Without the long-term phase drift, the short-term stability of phase measurement precision 16.2 mrad is shown as a gray area. Through noise analysis, we identified the inconsistency of SPADs and shot noise as the primary sources of noise. The shot-noise-induced phase noise was calculated to be 4.3 mrad. The inconsistency of SPADs would inevitably introduce phase noise under various free-space channel attenuation, estimated as 15.6 mrad. More details are provided in the Methods.